20. The specific rotation of (S)-alanine is +8.5°. If a mixture of alanine enantiomers is 80% S and 20% R, what is the specific rotation of mixture?

A) +14.17 B) -14.17 C) +10.62 D) +42.5
Answer #1

Specific rotation = + 8.50 enantiomeric excess = 7.2-1. S = 20% 804. 5-60% el = -60% Now, nuxture = specific rotation - ce xhope this helps...please rate if this is helpful..thank you